# Waveform Design of Multi-User-Multi-Target ISAC System based on Kullback–Leibler Divergence

### v1

@@ III-C Conditional KLD for an Arbitrary Precoding Matrix $\mathbf{W}_{\mathrm{c}}$ and Given Channel Matrix
A conditional KLD for communication, in terms of the bamforming matrix $\mathbf{W}_{\mathrm{c}}$ and for a given channel matrix $\mathbf{H}$, is required to formulate an optimization problem for which $\mathbf{W}_{\mathrm{c}}$ can be designed based on the channel side information matrix. Accordingly, by considering the received signal representation in (2), the conditional mean vectors and variance can be obtained as follows,
@@
where the assumptions of zero-mean, temporal whiteness, and wide-sense stationarity of the communication and radar signals, as well as their mutual orthogonality, have been utilized, it is important noting that the variance of the received signal for a given pair of symbols $\{s_{c,k,l}^{(n)},s_{c,k,l}^{(m)}\}\forall\{m,n\},m\neq n$ is equal. Therefore, the covariance is,
@@
Substituting the mean vectors and covariance matrices from (22), and (24) into the KLD expression given in (7), we can obtain the KLD for the $k$-th UE in communication subsystem for each possible pair of unequal data symbols $\{s_{\mathrm{c},k,l}^{(n)},s_{\mathrm{c},k,l}^{(m)}\}$, $n\neq m$, this is done as follows,
$\mathrm{KLD}_{k,m\rightarrow n}=\frac{d_{\mathrm{c},k}^{-\zeta}}{2\ln 2}\Bigg{(}\!\!\left(\mathbf{h}_{k}^{T}\mathbf{w}_{\mathrm{c},k}(s_{\mathrm{c},k,l}^{(m)}-s_{\mathrm{c},k,l}^{(n)})\right)^{H}\times$
$\Bigg{(}d_{\mathrm{c},k}^{-\zeta}\!\!\!\sum_{i=1,i\neq k}^{K}\!\!\!\|\mathbf{h}_{k}^{T}\mathbf{w}_{\mathrm{c},i}\|^{2}+\sigma_{\eta}^{2}\Bigg{)}^{\!\!-1}\!\!\!\!\!\!\left(\mathbf{h}_{k}^{T}\mathbf{w}_{\mathrm{c},k}(s_{\mathrm{c},k,l}^{(m)}-s_{\mathrm{c},k,l}^{(n)})\right)\!\!\Bigg{)},$ (25)
@@ IV-A Composite detection and response matrix estimation
The main objective for typical radar systems is to detect targets and estimate the matrix response associated with each one. After separating different signals associated with different targets, such problem can be formulated for each target as a composite detection-estimation problem as follows,
@@
This joint optimization problem aims to simultaneously detect the presence of a target (determining ^$\hat{\mathcal{H}}_{q}$) and estimate its response matrix ($\hat{\mathbf{H}}_{t}$) if present. Now, considering the hypotheses $\mathcal{H}_{0}$ and $\mathcal{H}_{1}$, the PDF of the received signals vector $\mathbf{y}_{\mathrm{r},t,l}$ under each of these hypotheses can be shown as a multivariate normal density, which can be written as
@@
where $\mathbf{R}_{1,t}=\mathbf{H}_{t}\mathbf{R}_{t,l}\mathbf{H}_{t}^{H}+\sigma_{n}^{2}\mathbf{I}_{N}$, and $\mathbf{R}_{t,l}=\mathbf{x}_{t,l}\>\mathbf{x}_{t,l}^{H}\>d_{\mathrm{r},t}^{\zeta}$. The generalized likelihood ratio test (GLRT) for the $t$-th target is used to determine target existence or absence by comparing the likelihoods under the two hypotheses. The GLRT is derived as follows,
@@
However, the PDF does not depend on $\mathbf{H}_{t}$ under $\mathcal{H}_{0}$. Therefore, the target response matrix is typically estimated under $\mathcal{H}_{1}$, thus by using the maximum likelihood estimation theorem we obtain
@@
Since the noise $\mathbf{n}_{t,l}$ is Gaussian and the $\mathbf{H}_{t}$ is normally distributed, the likelihood function is Gaussian, and the maximization leads to the least squares solution. Therefore, this maximization reduces to the following form,
@@
This least squares solution provides an unbiased estimate of $\mathbf{H}_{t}$. Substituting the PDF in (28) into the GLRT, we obtain the following,
@@
where $\hat{\mathbf{R}}_{1,t}=\hat{\mathbf{H}}_{t}\mathbf{R}_{t,l}\hat{\mathbf{H}}_{t}^{H}+\sigma_{n}^{2}\mathbf{I}_{N}$ and $\tau_{0}$ is the detection threshold that is a design parameter, for example, according to Neyman-Pearson lemma it can be determined to obtain a certain false alarm rate. Taking the logarithm and rearranging terms, we obtain the following expression, which now clearly includes the estimated target response matrix $\hat{\mathbf{H}}_{t}$,
@@
where $\tau_{1}=\ln\tau_{0}+N\ln\sigma_{n}^{2}-\ln\det(\hat{\mathbf{R}}_{1,t})$. This final expression correctly reflects the comparison between the hypotheses $\mathcal{H}_{0}$ and $\mathcal{H}_{1}$, incorporating the estimated target response matrix $\hat{\mathbf{H}}_{t}$ in the covariance matrix $\mathbf{R}_{1,t}$. In order to utilize the whole $L$ snapshots whose received signals are independent, we average as follows,
@@
where $\tau_{2}$ is the detection threshold that accounts for all snapshots and could be designed according to the Neyman-Pearson lemma as illustrated above.
@@ IV-B The KLD of radar system
Leveraging this asymptotic property and utilizing the $L$ number of snapshots, we can now derive the $\mathrm{KLD}_{\mathrm{r},t}^{\mathcal{H}_{0}\rightarrow\mathcal{H}_{1}}$ for the radar subsystem using the probability density functions established in (28), and (7) as follows,
@@
where $\mathbf{R}_{2,t}=\frac{1}{L}\sum_{l=1}^{L}\mathbf{R}_{1,t}=\mathbf{H}_{t}\frac{1}{L}\sum_{l=1}^{L}\mathbf{R}_{t,l}\mathbf{H}_{t}^{H}+\sigma_{n}^{2}\mathbf{I}_{N}=\mathbf{H}_{t}\mathbf{R}_{t}\mathbf{H}_{t}^{H}+\sigma_{n}^{2}\mathbf{I}_{N}$, and $\mathrm{det}(.)$ is the determinant operation. It should be noted that $\mathrm{KLD}_{\mathrm{r},t}=\frac{1}{2}\left(\mathrm{KLD}_{\mathrm{r},t}^{\mathcal{H}_{0}\rightarrow\mathcal{H}_{1}}+\mathrm{KLD}_{\mathrm{r},t}^{\mathcal{H}_{1}\rightarrow\mathcal{H}_{0}}\right)$, where in this case $\mathrm{KLD}_{\mathrm{r},t}^{\mathcal{H}_{0}\rightarrow\mathcal{H}_{1}}=\mathrm{KLD}_{\mathrm{r},t}^{\mathcal{H}_{1}\rightarrow\mathcal{H}_{0}}$.
@@ V Radar Waveform Optimisation
Here, the radar waveform is optimised while using the ZF beamforming scheme for the communication subsystem. The optimisation problem can be shown as follows,
@@
where $\mathrm{KLD}_{\mathrm{r},t|\mathbf{W}_{\mathrm{r},t},\mathbf{H}_{t}}$ is defined in (36). After substituting (36) in (37a), can be written as $\mathcal{P}_{2}$ that is given on page V. This optimization problem focuses on the radar waveform while considering the whole ISAC system performance. The constraint on radar power (37d) is explicit because the radar waveform is our primary optimization variable. The communication power is implicitly considered through the ZF beamforming scheme and the constraint (37c) on communication performance. While the communication power is not directly optimized, it is indirectly influenced by the radar power allocation. Specifically, the communication subsystem utilizes its allocated power through the ZF beamforming scheme to distribute power to each UE, with weights predetermined based on the channel state information. This approach allows us to focus on the interplay between subsystems which is mainly governed by waveform design and interference management while maintaining system-wide performance.
@@
To Tackle the optimisation problem in (38), we utilize the Projected Gradient method with a penalty function [36]. This method is chosen for its suitability to the non-linear, non-convex optimization problem [37]. This approach incorporates complex KLD constraints via the penalty function, while efficiently handles the radar power constraint through projection. It offers a balance between the solution quality and the computational efficiency, crucial for potential real-time ISAC applications. The method’s scalability with problem dimensionality further justifies the selection of this method for the radar waveform optimization task in the ISAC setup illustrated in the problem formulation. The optimization variable $\mathbf{\dot{W}}_{\mathrm{r}}\in\mathbb{C}^{N\times L\times T}$ is a three-dimensional tensor, where $\mathbf{w}_{\mathrm{r},t,l}$, $\mathbf{W}_{\mathrm{r},t}$, and $\mathbf{W}_{\mathrm{r},l}$ are all parts of this tensor. We define the objective function $f(\mathbf{\dot{W}}_{\mathrm{r}})$ and a penalty function $p(\mathbf{\dot{W}}_{\mathrm{r}})$ to handle the constraints in (37b) and (37c),
@@
where $\mathbf{G}^{(n)}=\nabla f(\mathbf{\dot{W}}_{\mathrm{r}}^{(n)})-\rho_{n}\nabla p(\mathbf{\dot{W}}_{\mathrm{r}}^{(n)})$ is the gradient direction, combining the objective function gradient and the penalty function gradient, $\rho_{n}$ is the penalty parameter, which balances the importance of the objective function and the constraint satisfaction, $\alpha_{n}$ is the step size, and $\mathcal{P}$ is the projection onto the feasible set defined by the power constraint, ensuring that the power constraint in (37d) is always satisfied. The projection mechanism is shown in the following equation,
@@
where $c\in(0,1)$. This condition, known as the Armijo condition, ensures a sufficient decrease in the penalized objective function [38]. The step size is then set as $\alpha_{n}=\beta^{m}\alpha_{n-1}$, where $\beta\in(0,1)$ is a reduction factor. The penalty parameter $\rho_{n}$ is updated as follows,
@@
Algorithm 1 Projected Gradient Method with Penalty for Radar Waveform Optimization
@@
Algorithm 1 implements the projected gradient method with penalty for the radar waveform optimization problem. It iteratively updates the waveform $\mathbf{\dot{W}}_{\mathrm{r}}$ by moving in the direction of the total gradient $\mathbf{G}^{(n)}$, which combines the gradients of the objective function gradient and the penalty function. The projection operator $\mathcal{P}$ ensures that the power constraint is always satisfied. This projection is crucial as it maintains the feasibility of the solution with respect to the power constraint throughout the optimization process. The backtracking line search procedure is used to determine an appropriate step size $\alpha_{n}$ at each iteration. This adaptive step size selection helps to ensure a sufficient increase in the objective function while maintaining the stability of the algorithm. The line search condition can be expressed as,
@@
where $\phi(\mathbf{\dot{W}}_{\mathrm{r}})=f(\mathbf{\dot{W}}_{\mathrm{r}})-\rho_{n}p(\mathbf{\dot{W}}_{\mathrm{r}})$ is the penalized objective function. The penalty parameter $\rho_{n}$ is updated adaptively based on the constraint violation. This adaptive scheme allows the algorithm to balance between optimizing the objective function and satisfying the constraints. As the optimization progresses, if constraints are consistently violated, the increasing penalty parameter will put more emphasis on constraint satisfaction.
@@ VI Communication optimisation
The communication subsystem optimization employs the gradient-assisted interior point method [39], contrasting with the projected gradient method with penalty used for radar optimization, due to the fundamental difference in optimization variable structure. While radar optimization involves a three-dimensional tensor $\mathbf{\dot{W}}_{\mathrm{r}}\in\mathbb{C}^{N\times L\times T}$, communication beamforming optimizes over a two-dimensional matrix $\mathbf{W}_{c}\in\mathbb{C}^{N\times K}$. This reduced dimensionality makes IPM particularly suitable, potentially offering faster convergence and more precise solutions for the matrix optimization problem. The communication waveform is optimised while using conventional identity covariance matrix design for the radar subsystem with a covariance matrix set to $\mathbf{R}_{w}=\mathbf{I}_{N\times N}\quad\forall t\in T$. The optimisation problem can be shown as follows,
@@
It is important to note that the radar power is uniformly distributed among the potential targets, as the radar covariance matrix ensures an omnidirectional beam flow. To solve this optimization problem, we need to derive the gradient of the objective function and the constraints. Let’s define the objective function as,
@@
Now, to compute the gradient of the objective function, we need to derive the gradients of each objective function $\mathbf{W}_{\mathrm{c}}$ components.Let $a_{k}=d_{\mathrm{c},k}^{\zeta}|\mathbf{h}_{k}^{T}\mathbf{w}_{\mathrm{c},k}|^{2}$. Using the chain rule and the properties of matrix derivatives, the gradients can be expressed as,
@@
Next, we need to consider the constraints of our optimization problem. The power constraint is given by,
@@
The gradient of this constraint is derived straightforwardly as follows,
@@
Algorithm 2 Interior Point Method for Communication Optimization
@@
Solving the optimization problem defined in equations (50a)-(50d) involves employing the Interior Point Method (IPM) as outlined in Algorithm 2. This method efficiently handles constrained optimization using key parameters: an initial barrier parameter $\tilde{\mu_{0}}$, a barrier reduction factor $\tilde{\gamma}$, and a convergence tolerance $\epsilon$. The algorithm iteratively computes the objective function, constraints, and their gradients, checks the Karush-Kuhn-Tucker (KKT) conditions, and updates the solution using a Newton system and line search. The barrier parameter $\tilde{\mu}$ is reduced as $\tilde{\mu}=\tilde{\gamma}\tilde{\mu}$ in each iteration, gradually enforcing constraints more strictly. This process continues until $|\mathbf{W}_{c}^{(n+1)}-\mathbf{W}_{c}^{(n)}|_{F}<\epsilon$ or the maximum iterations are reached, yielding the optimal beamforming matrix $\mathbf{W}_{c}^{*}$ that maximizes average communication KLD while satisfying radar and power constraints.

### v2

@@ III-C Conditional KLD for an Arbitrary Precoding Matrix $\mathbf{W}_{\mathrm{c}}$ and Given Channel Matrix
A conditional KLD for communication, in terms of the beamforming matrix $\mathbf{W}_{\mathrm{c}}$ and for a given channel matrix $\mathbf{H}$, is required to formulate an optimisation problem. Considering the received signal representation in (2), the conditional mean vectors and variance are,
@@
Note that, the variance of the received signal for a given pair of symbols $\{s_{c,k,l}^{(n)},s_{c,k,l}^{(m)}\}\forall\{m,n\},m\neq n$ is equal. Thus, the covariance is,
@@
Substituting the mean vectors and covariance matrices from (22), and (24) into the KLD expression given in (7), we can obtain the KLD for the $k$-th UE in the communication subsystem for each possible pair of unequal data symbols $\{s_{\mathrm{c},k,l}^{(n)},s_{\mathrm{c},k,l}^{(m)}\}$, $n\neq m$, this is done as follows,
$\mathrm{KLD}_{k,m\rightarrow n}=\frac{d_{\mathrm{c},k}^{-\zeta}}{2\ln 2}\Bigg{(}\!\!\left(\mathbf{h}_{k}^{T}\mathbf{w}_{\mathrm{c},k}(s_{\mathrm{c},k,l}^{(m)}-s_{\mathrm{c},k,l}^{(n)})\right)^{H}$
×$\times\Bigg{(}d_{\mathrm{c},k}^{-\zeta}\!\!\!\sum_{i=1,i\neq k}^{K}\!\!\!\|\mathbf{h}_{k}^{T}\mathbf{w}_{\mathrm{c},i}\|^{2}+\sigma_{\eta}^{2}\Bigg{)}^{\!\!-1}\!\!\!\!\!\!\left(\mathbf{h}_{k}^{T}\mathbf{w}_{\mathrm{c},k}(s_{\mathrm{c},k,l}^{(m)}-s_{\mathrm{c},k,l}^{(n)})\right)\!\!\Bigg{)},$ (25)
@@ IV-A Composite detection and response matrix estimation
The main objective of typical radar systems is to detect targets and estimate the matrix response for each one. After separating signals associated with different targets, we formulate a composite detection-estimation problem as follows,
@@
This joint optimisation problem aims to simultaneously detect the presence of a target (determining ^$\hat{\mathcal{H}}_{q}$) and estimate its response matrix ($\hat{\mathbf{H}}_{t}$) if present. For the hypotheses $\mathcal{H}_{0}$ and $\mathcal{H}_{1}$, the PDF of the received signals vector $\mathbf{y}_{\mathrm{r},t,l}$ under each hypothesis can be shown as a multivariate normal density,
@@
where $\mathbf{R}_{1,t}=\mathbf{H}_{t}\mathbf{R}_{t,l}\mathbf{H}_{t}^{H}+\sigma_{n}^{2}\mathbf{I}_{N}$, and $\mathbf{R}_{t,l}=\mathbf{x}_{t,l}\>\mathbf{x}_{t,l}^{H}\>d_{\mathrm{r},t}^{\zeta}$. The generalised likelihood ratio test (GLRT) for the $t$-th target is used to determine target existence or absence by comparing the likelihoods under the two hypotheses, derived as,
@@
As the PDF does not depend on $\mathbf{H}_{t}$ under $\mathcal{H}_{0}$, the target response matrix is estimated under $\mathcal{H}_{1}$, thus by using the maximum likelihood estimation theorem we obtain,
@@
The likelihood function is Gaussian, and the maximisation leads to the least squares solution,
@@
This least squares solution provides an unbiased estimate of $\mathbf{H}_{t}$. Substituting the PDF in (28) into the GLRT, we obtain,
@@
where $\hat{\mathbf{R}}_{1,t}=\hat{\mathbf{H}}_{t}\mathbf{R}_{t,l}\hat{\mathbf{H}}_{t}^{H}+\sigma_{n}^{2}\mathbf{I}_{N}$ and $\tau_{0}$ is the detection threshold. Taking the logarithm and rearranging, we obtain,
@@
where $\tau_{1}=\ln\tau_{0}+N\ln\sigma_{n}^{2}-\ln\det(\hat{\mathbf{R}}_{1,t})$. This expression compares hypotheses $\mathcal{H}_{0}$ and $\mathcal{H}_{1}$, incorporating the estimated target response matrix $\hat{\mathbf{H}}_{t}$ in the covariance matrix $\mathbf{R}_{1,t}$. To use all $L$ snapshots with independent signals, we average as,
@@
where $\tau_{2}$ is the detection threshold considering all snapshots and can be designed per the Neyman-Pearson lemma.
@@ IV-B The KLD of radar system
Using this property, we can now derive the $\mathrm{KLD}_{\mathrm{r},t}^{\mathcal{H}_{0}\rightarrow\mathcal{H}_{1}}$ for the radar subsystem using the probability density functions established in (28), and (7) as follows,
@@
where $\mathbf{R}_{2,t}=\frac{1}{L}\sum_{l=1}^{L}\mathbf{R}_{1,t}=\mathbf{H}_{t}\frac{1}{L}\sum_{l=1}^{L}\mathbf{R}_{t,l}\mathbf{H}_{t}^{H}+\sigma_{n}^{2}\mathbf{I}_{N}=\mathbf{H}_{t}\mathbf{R}_{t}\mathbf{H}_{t}^{H}+\sigma_{n}^{2}\mathbf{I}_{N}$, and $\mathrm{det}(.)$ is the determinant operation. $\mathrm{KLD}_{\mathrm{r},t}=\frac{1}{2}\left(\mathrm{KLD}_{\mathrm{r},t}^{\mathcal{H}_{0}\rightarrow\mathcal{H}_{1}}+\mathrm{KLD}_{\mathrm{r},t}^{\mathcal{H}_{1}\rightarrow\mathcal{H}_{0}}\right)$, where in this case $\mathrm{KLD}_{\mathrm{r},t}^{\mathcal{H}_{0}\rightarrow\mathcal{H}_{1}}=\mathrm{KLD}_{\mathrm{r},t}^{\mathcal{H}_{1}\rightarrow\mathcal{H}_{0}}$.
@@ V Radar Waveform Optimisation
In this section, the radar waveform is optimised while using the ZF beamforming scheme for the communication subsystem. The optimisation problem is,
@@
where $\mathrm{KLD}_{\mathrm{r},t|\mathbf{W}_{\mathrm{r},t},\mathbf{H}_{t}}$ is defined in (36). After substituting (36) in (37a), can be written as $\mathcal{P}_{2}$ that is given on page V. This optimisation problem focuses on the radar waveform while considering the whole ISAC system performance. The constraint on radar power (37d) is explicit because the radar waveform is our primary optimisation variable. The communication power is implicitly considered through the ZF beamforming scheme and the constraint (37c) on communication performance. Specifically, the communication subsystem utilises its allocated power through the ZF beamforming scheme to distribute power to each UE, with weights predetermined based on the channel state information.
@@
To solve the optimisation problem in (38), we utilise the Projected Gradient method with a penalty function [36]. This method is chosen for its suitability to the non-linear, non-convex optimisation problem [37]. This method uses a penalty function for KLD constraints and handles radar power via projection, balancing the solution quality and computational efficiency for real-time ISAC applications. Its scalability makes it suitable for radar waveform optimisation in ISAC. The optimisation variable $\mathbf{\dot{W}}_{\mathrm{r}}\in\mathbb{C}^{N\times L\times T}$ is a three-dimensional tensor, where $\mathbf{w}_{\mathrm{r},t,l}$, $\mathbf{W}_{\mathrm{r},t}$, and $\mathbf{W}_{\mathrm{r},l}$ are all parts of this tensor. We define the objective function $f(\mathbf{\dot{W}}_{\mathrm{r}})$ and a penalty function $p(\mathbf{\dot{W}}_{\mathrm{r}})$ to handle the constraints in (37b) and (37c),
@@
where $\mathbf{G}^{(n)}=\nabla f(\mathbf{\dot{W}}_{\mathrm{r}}^{(n)})-\rho_{n}\nabla p(\mathbf{\dot{W}}_{\mathrm{r}}^{(n)})$ is the gradient direction, combining the objective function gradient and the penalty function gradient, $\rho_{n}$ is the penalty parameter, balancing the objective and constraint satisfaction, $\alpha_{n}$ is the step size, and $\mathcal{P}$ is the projection onto the feasible set defined by the power constraint, ensuring that the power constraint in (37d) is always satisfied. The projection mechanism is shown as,
@@
where $c\in(0,1)$. This condition, known as the Armijo condition, ensures a sufficient decrease in the penalised objective function [38]. The step size is set as $\alpha_{n}=\beta^{m}\alpha_{n-1}$, with $\beta\in(0,1)$ as a reduction factor. The penalty parameter $\rho_{n}$ is updated as follows,
@@
Algorithm 1 Projected Gradient Method with Penalty for Radar Waveform Optimisation
@@
The projection operator $\mathcal{P}$ ensures that the power constraint is always satisfied. The backtracking line search procedure is used to determine an appropriate step size $\alpha_{n}$ at each iteration. This adaptive step size selection helps to ensure a sufficient increase in the objective function while maintaining the stability of the algorithm. The line search condition can be expressed as,
@@
where $\phi(\mathbf{\dot{W}}_{\mathrm{r}})=f(\mathbf{\dot{W}}_{\mathrm{r}})-\rho_{n}p(\mathbf{\dot{W}}_{\mathrm{r}})$ is the penalised objective function. The penalty parameter $\rho_{n}$ is updated adaptively based on the constraint violation, allowing the algorithm to balance optimisation and constraint satisfaction. If constraints are consistently violated, the increasing penalty will emphasise constraint adherence.
@@ VI Communication optimisation
The communication subsystem optimisation uses the gradient-assisted IPM [39], in contrast to the projected gradient method with penalty used for radar optimisation, due to the different optimisation variable structures. Radar optimisation involves a three-dimensional tensor $\mathbf{\dot{W}}_{\mathrm{r}}\in\mathbb{C}^{N\times L\times T}$, while communication beamforming optimises a two-dimensional matrix $\mathbf{W}_{c}\in\mathbb{C}^{N\times K}$. This reduced dimensionality makes IPM particularly suitable, potentially offering faster convergence and more precise solutions. The communication waveform is optimised while using the conventional identity covariance matrix design for the radar subsystem, with $\mathbf{R}_{w}=\mathbf{I}_{N\times N}\quad\forall t\in T$. The optimisation problem can be expressed as follows,
@@
Note, that the radar power is uniformly distributed among potential targets, as the radar covariance matrix ensures omnidirectional beam flow. To solve this optimisation problem, we must derive the gradients of the objective function and the constraints. Let’s define the objective function as follows,
@@
To compute the gradient of the objective function, we need to derive the gradients of each component of $\mathbf{W}_{\mathrm{c}}$. Let $a_{k}=d_{\mathrm{c},k}^{\zeta}|\mathbf{h}_{k}^{T}\mathbf{w}_{\mathrm{c},k}|^{2}$. Using the chain rule and properties of matrix derivatives, the gradients can be expressed as follows,
@@
Next, we need to consider the constraints of our optimisation problem. The power constraint is given by,
@@
The gradient of this constraint is derived as follows,
@@
Algorithm 2 Interior Point Method for Communication Optimisation
@@
Solving the optimisation problem defined in equations (50a)-(50d) uses the IPM, as detailed in Algorithm 2. This method efficiently addresses constrained optimisation with key parameters: an initial barrier parameter $\tilde{\mu_{0}}$, a barrier reduction factor $\tilde{\gamma}$, and a convergence tolerance $\epsilon$. The algorithm iteratively computes the objective function, constraints, and their gradients, checks the Karush-Kuhn-Tucker (KKT) conditions and updates the solution using a Newton system and line search. The barrier parameter $\tilde{\mu}$ is updated as $\tilde{\mu}=\tilde{\gamma}\tilde{\mu}$ in each iteration, gradually enforcing constraints more strictly. This continues until $|\mathbf{W}_{c}^{(n+1)}-\mathbf{W}_{c}^{(n)}|_{F}<\epsilon$ or the maximum iterations are reached, yielding the optimal beamforming matrix $\mathbf{W}_{c}^{*}$ that maximises the communication KLD while satisfying radar and power constraints.